OpenStudy (anonymous):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8).

OpenStudy (anonymous):

@TheDrowzeeAbra

OpenStudy (anonymous):

First you need to find the slope. Do you know how to do that?

OpenStudy (anonymous):

im just checking my answers i juts wanna know if this is right?

OpenStudy (anonymous):

y = 1/3x + 29/3

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

The slope should be -1/3. Try to find the new b

OpenStudy (anonymous):

y = -1/3x - 29/3

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

I got y = -1/3 + 43/3

OpenStudy (anonymous):

Do you get how?